OpenStudy (anonymous):

the perimeter of a rectangle is 56m. The length is 1m more than twice the width. Find the dimensions. L= W=

OpenStudy (anonymous):

Hi there! Let's solve a linear system! Call the length L and the width W. First, can you write an equation that relates L and W to the perimeter?

OpenStudy (anonymous):

2L+2W=56

OpenStudy (anonymous):

Yep, you got it. Now, can you write another equation, also using L and W, that expresses " The length is 1m more than twice the width"?

OpenStudy (anonymous):

L=2W+1

OpenStudy (anonymous):

Perfect!! Now solve the system using the substitution method. You know L=2W+1, so substitute "2W+1" in every place L appears in the first equation.

OpenStudy (anonymous):

2(2W+1)+2W=56

OpenStudy (anonymous):

Perfect. Now expand that first part out by distributing the 2. Then collect like terms.

OpenStudy (anonymous):

4W+2+2W=56

OpenStudy (anonymous):

right on. Now collect all the W terms on the left and the constant terms on the right.

OpenStudy (anonymous):

2W=56

OpenStudy (anonymous):

Oops, not quite. You had 4W + 2 + 2W = 56, that was right. What do you get when you collect the W terms together? i.e. 4W + 2W?

OpenStudy (anonymous):

4W=56

OpenStudy (anonymous):

Let me write it again in a way that makes it look a little easier: 4W + 2 + 2W = 56 is the same as 4W + 2W + 2 = 56 Can you combine the 4W and 2W together?

OpenStudy (anonymous):

Pretend W is actually the word "marble". Add 4 marbles and 2 marbles.

OpenStudy (anonymous):

ok now you just lost me

OpenStudy (anonymous):

You understand how we got to 4W + 2W + 2 = 56, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

W=9

OpenStudy (anonymous):

OK, so now we combine the "W" terms together -- that is, the terms that have a W in them.

OpenStudy (anonymous):

Oops, yep, you skipped ahead - that's great! w = 9. Now substitute w=9 into either of the two equations you started with in order to find L.

OpenStudy (anonymous):

L=2(9)+1

OpenStudy (anonymous):

Now just solve for L.

OpenStudy (anonymous):

L=19

OpenStudy (anonymous):

you got it. you have the solution!